A certain chemical reaction releases 22.2kJ/g of heat for each gram of reactant consumed. How can you calculate the heat produced by the consumption of 1.5kg of reactant?
1
Expert's answer
2020-08-26T06:35:44-0400

The mass of the reactant is 1.5 kg or 1500 g. It should be multiplied by specific heat (heat for a gram):

Q=m×c=1500×22.2=33300(kJ).